Please explain why the answer is 0.588 m.
A double slit with a slit separation distance of 2
Answer #1

For this case we use the small angle approximation:

y_m = \frac{(m+1/2)\lambda L}{d}

The fourth dark fringe is for m = 3

y= \frac{(3+1/2)(560 \times 10^{-9})(6.00)}{2\times 10^{-5}}= 0.588 m